Logiweb diagnose of kvanti

Up Help

Error in proof of JoinConjuncts(2conditions) Lemma expected MP2

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-12-15.UTC:00:19:10.164930 = MJD-54084.TAI:00:19:43.164930 = LGT-4672858783164930e-6